OpenStudy (phebe):

12. Using the discriminant, how many real number solutions does this equation have? 3x2 – 2 = 5x

OpenStudy (solomonzelman):

1) subtract 5x from both sides 2) set the equation in order ax^2+bx+c=0 3) plug in numbers inside the discriminant, which is b²-4(a)(c)

OpenStudy (phebe):

okkk

OpenStudy (phebe):

i did step 3

OpenStudy (solomonzelman):

and what do you get ?

OpenStudy (phebe):

wait a sec im wrong

OpenStudy (phebe):

wait the problem was 3x^2-2=5x

OpenStudy (solomonzelman):

yes, and when you subtract 5x from both sides you get \(\LARGE\color{blue}{ \bf 3x^2-5x-2=0 }\)

OpenStudy (phebe):

yess i did that too

OpenStudy (phebe):

so then do i solve from there

OpenStudy (phebe):

@SolomonZelman

OpenStudy (solomonzelman):

I am sure you have been solving quadratic equations using quadratic formulas before, haven't you ?

OpenStudy (dan815):

yep, or good old factoring

OpenStudy (phebe):

yess

OpenStudy (solomonzelman):

no... he needs the discriminant.

OpenStudy (solomonzelman):

or she...

OpenStudy (phebe):

im a girl lol

OpenStudy (dan815):